A body is thrown vertically upward with initial velocity u reaches a maximum height in 6 seconds the ratio of distances travelled by the body in the first second and 7thsecond is

Grade:9

1 Answers

Arun
4 years ago
 
for full motion,
v=u-gt
=>0=u-gt
=>u=gt=10*6=60 ms^-1
 
dist travelled in first second=s1= u-(g/2)*(2n-1) = 60-5*1=55
dist travelled in seventh second=s7= u-(g/2)*(2n-1) =60-5*13=-5 (negative as direction is different)
so ratio =s1:|s7|
=55:5
=11:1